OpenStudy (anonymous):

How would you go about solving: (arctan(1/x))' ?

OpenStudy (anonymous):

Without using that (arctan(u))'=1/(1+u^2).

OpenStudy (dls):

can u rewrite in equation form? what is " ' " ?

OpenStudy (anonymous):

Derivative.

OpenStudy (anonymous):

so your real question is how we find (arctan(u))'=1/(1+u^2) ? cause when we know it we can use it

OpenStudy (anonymous):

Basically. :D

OpenStudy (anonymous):

y = arctan(x) we want to find y' now: tan(y) = x so take the derivative with respect to x sec^2(y) * y' = 1 y' = cos^2(y) y' = cos^2(arctan(x))

OpenStudy (anonymous):



OpenStudy (anonymous):

I never worked with the function sec. :/

OpenStudy (anonymous):

sec = 1/cos

OpenStudy (anonymous):

y' = cos^2(arctan(x)) according to my drawing : arctan(x) = t so y' = cos^2(t) now according to the drawing: cos(t) = 1/sqrt(1+x^2) so cos^2(t) = 1/(1+x^2) so y' = 1/(1+x^2)

OpenStudy (anonymous):

"I never worked with the function sec. :/" so you know (tan(x))' = 1/(cos^2(x)) i just use sec(x) = 1/cos(x) so (tan(x))' = sec^2(x)

OpenStudy (anonymous):

That's a real cool alternative approach without using limits, what is the domain of the relation sec(x)=1/cos(x)?

OpenStudy (anonymous):

this is just a definition.. instead of writing 1/cos(x) we write sec(x)

OpenStudy (anonymous):

Yes I know, but what is the domain?

OpenStudy (anonymous):

Is it applicable theoretically even though it's derived from a geometric relation?

OpenStudy (anonymous):

look at the function : http://www.wolframalpha.com/input/?i=secx&dataset=

OpenStudy (anonymous):

there is also csc which is 1/sin

OpenStudy (anonymous):

"Is it applicable theoretically even though it's derived from a geometric relation?" you ask if the derivative of arctan ? yes .. applicable

OpenStudy (anonymous):

So both those relations above have the domain C?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

I mean if the relation is applicable for all the possible values of t and x (even though your figure above might not even be a triangle). The relation being sec(x)=1/cos(x)

OpenStudy (anonymous):

Nevermind, I was just being curious about the relation sec(x)=1/(cos(x)); it was very beautiful.

OpenStudy (anonymous):

the relation sec(x)=1/cos(x) is independent of the triangle.

OpenStudy (anonymous):

Okay. :) Do you know a proof of this?

OpenStudy (anonymous):

Nevermind, I was just being curious about the relation sec(x)=1/(cos(x)); it was very beautiful.

OpenStudy (anonymous):

Okay. :) Do you know a proof of this?

OpenStudy (anonymous):

the triangle was there to help us find what is cos^2(arctan(x)) in terms of x.

OpenStudy (anonymous):

Oooh.

OpenStudy (anonymous):

we dont need any proof for sec(x) = 1/(cos(x)) this is a definition .. i can make a new one : coolsector(x) = x^2 + 6

OpenStudy (anonymous):

sec(x) is just a name for the function 1/cos(x)

OpenStudy (anonymous):

I see! I got it mixed up. Thanks :)

OpenStudy (anonymous):

yw :)